Question 1 – M&A Blended Offer

Sprint is planning on acquiring Nextel. The situation for both firms before the transaction is as outlined below:

Sprint before the transaction:

  • 1,400 million shares outstanding at a market price of $25 per share
  • Market value of debt is $5,000 million
  • No excess cash

Nextel before the transaction:

  • 1,030 million shares outstanding at a market price of $30 per share
  • Market value of debt is $5,000 million
  • No excess cash

Transaction details:

  • Sprint will pay $2 per share of Nextel and will also exchange each share of Nextel for 1.1661 shares of Sprint
  • Sprint will finance the cash component of the offer by taking on an acquisition loan
  • Sprint will assume the outstanding debt of Nextel
  • Synergies from the acquisition will be $12,000 million (including any tax shields from the acquisition loan)
  1. What is the enterprise value of Sprint after its acquisition of Nextel (???????+??????)?
  2. What is the market value of equity of Sprint after its acquisition of Nextel (???????+??????)?
  3. What is the share price of Sprint after its acquisition of Nextel (???????+??????)?
  4. What is the Price Paid by Sprint for the acquisition of Nextel?
  5. How much value did Sprint create for its shareholders through the acquisition?

Question 2 – IPO Underwriter Spreads and Money Left on the Table

Below is a table with data for three companies that conducted and IPO. All of these IPOs were firm commitment IPOs underwritten by the same underwriter.

Firm

Offer Price Per Share (in $)

First Closing Market Price Per Share (in $)

# Shares offered (in $mn)

Underwriter Spread (in %)

A

$38.00

$95.06

23

7.0%

B

$33.00

$34.19

478

2.0%

C

$15.00

$23.43

58

5.0%

  1. For each of the three companies, calculate the $-amount (in $mn) of fees and discounts that the underwriter charged for the IPO.
  2. For each of the three companies, calculate how much money was left on the table (in $mn) in their IPOs.

Question 3 – IPO Winner’s Curse

Suppose DCB Inc. currently has 100mn shares outstanding and is attempting to sell 12mn new shares through an IPO. The underwriter believes if the IPO is successful, the true value of DCB’s equity will be either $1,120mn (prob. 40%) or $1,792mn (prob. 60%).

Suppose there are uninformed investors who would be willing to buy as much as 12 million shares, provided that their expected return is non-negative. Uninformed investors don’t know the true value of the firm, they only know the probability distribution of the potential firm values.

Additionally, there are informed investors, who know the true value of the firm, and are willing to submit bids for up to 4 million shares, provided that the offer price is lower than the true value.

Assume that the underwriter will assign shares on a pro-rata basis if the offer is oversubscribed.

  1. What would the fair IPO offer price be if all investors are uninformed?
  2. How does the underwriter need to set the IPO offer price for the offer to clear?
Answer #1

You have asked multiple unrelated questions in the same post. Further each question has multiple sub parts. I have addressed all the sub parts of the first question. Please post the balance questions separately.

What is the enterprise value of Sprint after its acquisition of Nextel (???????+??????)?

Sprint will pay $2 per share of Nextel and will also exchange each share of Nextel for 1.1661 shares of Sprint

Hence, value per share of Nextel = 2 + 1.1661 x Price per share of Sprint = 2 + 1.1661 x 25 = $ 31.15

EV of Sprint = Equity value + existing debt + new debt - cash consideration paid towards acquisition of Nextel

Since, new debt = cash consideration paid towards acquisition of Nextel, hence EV of Sprint = Equity value + existing debt = 1,400 million shares x 25 per share + 5,000 = 40,000 million

EV of Nextel based on acquisition price = Equity value based on acquisition price + debt = 31.15 / share x 1,030 million shares + 5,000 =  37,087.08

Hence, EV of Sprint after its acquisition of Nextel (???????+??????) = EV of Sprint + EV of Nextel + Synergies = 40,000 +  37,087.08 + 12,000 = $  89,087.08 million

What is the market value of equity of Sprint after its acquisition of Nextel (???????+??????)?

Market value of equity = Enterprise value - debt =  89,087.08 - existing debt of ESprint - existing debt of Nextel - fresh debt for acquisition =  89,087.08 - 5,000 - 5,000 - 2 x 1,030 = $ 77,027.08 million

What is the share price of Sprint after its acquisition of Nextel (???????+??????)?

Number of shares outstanding after acquisition = old number of shares + new shares issued = 1,400 + 1.1661 x 1,030 =  2,601.08 million

Hence, per share price = Equity value / Number of shares outstanding =  77,027.08 /  2,601.08 = $  29.61 per share

What is the Price Paid by Sprint for the acquisition of Nextel?

Sprint will pay $2 per share of Nextel and will also exchange each share of Nextel for 1.1661 shares of Sprint

Hence, price paid per share of Nextel = 2 + 1.1661 x Price per share of Sprint = 2 + 1.1661 x 25 = $ 31.15

Total consideration (Price) paid by Sprint for the acquisition of Nextel = 31.15 x 1,030 = $  32,087.08 million

How much value did Sprint create for its shareholders through the acquisition?

Value created = (New share price - old share price ) x Number of old shares outstanding = (29.61 - 25) x 1,400 = $  6,458.85 million